In a transformer, number of turns in the primary coil are 140 and that in the secondary coil are 280. If current in primary coil is 4 A, then that in the secondary coil is
(a) 4 A
(b) 2 A
(c) 6 A
(d) 10 A.

Grade:6

2 Answers

Navjyot Kalra
9 years ago

(b) Np = 140 , Ns = 280, Ip = 4A, Is = ?

For a transformer Is / Ip =NP / Ns

? Is /4 = 140 / 250 ? Is = 2A

Vasantha Kumari
9 years ago

We know that for a transformer I1N1 = I2N2 where I1 and I2 are the currents through the coil and N1 = N2 are the number of turns in the primary and secondary coils.

Therefore, I2 = I1N1/ N2

=4 x 140 / 280

= 2A

Thanks & Regards,

Vasantha Sivaraj,

askIITians faculty